# Quantum entanglement, coherence and Dark Matter

### v1

@@ Abstract
In this paper we consider the effect of quantum entanglement and coherence on the radiated intensity from a gas and its absorption capacity at thermal equilibrium or, more generally, under conditions where no population inversion exists. As was shown by Dicke (1954), although entangled states and coherence can lead to superradiance for specific modes of radiation, they can also bring subradiance through significant energy trapping in slow and dark states. While a finite separation between the atoms composing the gas will cause leaking of the trapped energy, we show how the combination of thermal equilibrium and quantum coherence mitigates this effect and leads to significantly reduced radiation intensity from the gas, rendering it dark and collision-less. Furthermore, we show how under the same conditions absorption of a radiation field incident on the gas can lead to higher attenuation levels than those predicted with Beer’s law. Beer’s law is recovered in the limit of complete decoherence. We apply our analysis to the atomic hydrogen 21 cm line and, considering the gas densities expected in Dark Matter halos, we find that quantum entanglement and coherence can potentially account for some of the Dark Matter known exist in these environments.
I Introduction
It has been known since Dicke’s original treatment of the spontaneous emission of coherent radiation [1] that the interaction between atoms composing a gas  [^1] Although we focus on atoms in this paper, our analysis and discussions apply to molecular gases as well. through their common electromagnetic field leads to the formation of entangled quantum mechanical states, with significant implications for the spontaneous emission rates of photons emanating from the gas. As Dicke pointed out, atoms should not be considered independently; rather, the entire gas must be treated as a single quantum mechanical system. While the material covered in his paper was broad, touching on several phenomena involving coherence and entanglement, much of the early work it inspired focused on superradiance.
@@ I Introduction
Dicke’s 1954 paper [1] also explicitly discussed another coherent radiation process, which later became known as subradiance. In contrast to superradiance, subradiance describes the suppression of photon emission compared to the typical spontaneous emission process. Dicke analyzed the trapping of energy in slow and dark states and its consequences for the radiation intensity emanating from a gas. Subradiance was experimentally verified later than superradiance [16] and has only recently been observed in large clouds of cold atoms [17]. The latter experimental result is particularly relevant to us as we explore the possible existence of subradiant systems in astrophysical contexts.
@@
Our paper is structured as follows. In Sec. II, we review key material from Dicke’s 1954 paper [1], focusing on energy trapping. In Sec. III, we extend this analysis by introducing equilibrium conditions within the framework of the master equation. In Sec. IV, we apply the material developed in the previous sections to astrophysical contexts using the Maxwell-Bloch equations. Specifically, we focus on the 21cm line of atomic hydrogen to investigate the potential realization of subradiance in environments with densities consistent with those of galactic Dark Matter halos. Finally, we discuss our results in Sec. V, while detailed calculations are provided in three appendices at the end.
II Entangled quantum states and coherent behavior
In this section we revisit parts of Dicke’s work and focus on the trapping of energy within an atomic ensemble brought about by the existence of Dicke states of differing symmetries. Accordingly, we follow Dicke [1] and express the Hamiltonian for the gas with
$\hat{H}=\hat{H}_{0}+\hbar\omega_{0}\sum_{j}\hat{R}^{3}_{j}+\hat{V},$ (1)
where $\hat{H}_{0}$ accounts for the translational motions of the atoms’ center-of-mass and their interactions, the second term for their internal energy and $\hat{V}$ for dipole interactions with the electromagnetic field. For simplicity we approximate the atoms as two-level systems at resonance with each other (i.e., moving at the same velocity) with an internal energy difference $\hbar\omega_{0}$; the summation in equation (1) is on all atoms. That is, if |$\ket{a}$ and |$\ket{b}$ are the lower and upper atomic states, respectively, then for the $j^{\mathrm{th}}$ atom $\hat{R}^{3}_{j}=\frac{1}{2}\left(\left|b\middle\rangle\middle\langle b\right|-\left|a\middle\rangle\middle\langle a\right|\right)$.
For our problem the interaction term reduces to
$\Omega_{k}=\frac{2\mu\mathcal{E}_{0}}{\hbar c}\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{k}^{\perp}\right)$ (3)
for a magnetic dipole transition, where $\mu$ and $\bm{\epsilon}_{\mu}$ are the magnitude of the magnetic dipole moment and its associated unit vector (assumed real), while that for the radiation mode’s polarization is (also real and) denoted by $\bm{\epsilon}_{k}^{\perp}$ (i.e., $\bm{k}=k\bm{\epsilon}_{k}$ and $\bm{\epsilon}_{k}\cdot\bm{\epsilon}_{k}^{\perp}=0$). The one-photon electric field is
$\mathcal{E}_{0}=\sqrt{\frac{\hbar\omega_{0}}{2\epsilon_{0}\mathcal{V}}}$ (4)
with $\epsilon_{0}$ the permittivity of vacuum and $\mathcal{V}$ the volume of quantization. We focus on magnetic dipole transitions in view of our application to the atomic hydrogen 21 cm line in Sec. IV. For an electrical dipole transition one has to substitute $\mu\rightarrow dc$ and $\bm{\epsilon}_{\mu}\rightarrow\bm{\epsilon}_{d}$ in equation (3), with $d$ the magnitude of the electric dipole moment.
$\hat{R}^{\pm}_{k}=\sum_{j}\hat{R}^{\pm}_{j}e^{\pm i\bm{k}\cdot\mathbf{r}_{j}}$ (5)
with $\mathbf{r}_{j}$ the position of atom $j$. For the single-atom raising and lowering operators we have $\hat{R}^{+}_{j}=\left|b\middle\rangle\middle\langle a\right|$ and $\hat{R}^{-}_{j}=\left|a\middle\rangle\middle\langle b\right|$ while, as we assume resonance, we write $k=\left|\mathbf{k}\right|=\omega_{0}/c$.
II.1 The two-atom problem
Let us consider the problem of two identical two-level atoms located a distance $z_{0}$ from one another; Atom 1 at $-z_{0}/2$ and Atom 2 at $z_{0}/2$ on the $z$-axis. For this system
with $\theta$ the angle of k relative to the $z$-axis.
Given that single-atom transitions take place between the lower |$\Ket{a}$ and upper |$\Ket{b}$ levels of energies $-\hbar\omega_{0}/2$ and $\hbar\omega_{0}/2$, respectively, appropriate internal states for the two-atom problem consist of [1, 19] (see Appendix C)
|$\Ket{1,1}=\Ket{bb}$ (7)
@@ II.1 The two-atom problem
Following Dicke [1], the kets |$\Ket{r,m}$ are defined with the ‘cooperative’ number $r$ and ‘inversion’ number $m$, which are analogous to spin-1/2 quantum numbers, with $r\geq 0$ and $\left|m\right|\leq r$ (note that |$\Ket{r,m}_{\theta}=\Ket{r,m}$ when $\cos\theta=0$).
Given equations (7)-(10) it is straightforward to calculate radiative transition rates between the different states upon applying Fermi’s Golden Rule with equations (2), (3) and (6). When $kz_{0}\neq 0$ the selection rules are then found to be $\Delta r=0,\pm 1$ and $\Delta m=\pm 1$, and the spontaneous emission rates per unit solid angle
@@
The signatures of entanglement and coherence are displayed in equations (13)-(14). That is, the enhancement of the transitions rates by a factor of ‘2’ is due to the entangled nature of the states given in equations (8) and (10), while the disappearance of the argument in the cosine and sine functions stems from coherence between the two emitted photons (i.e., when $\cos\theta^{\prime}=\cos\theta$). We therefore find the existence of an angular correlation between successively emitted photons. More precisely, if the system starts in the |$\Ket{1,1}$ state, then after the emission of the first photon it will for sure end up in the symmetric |$\Ket{1,0}_{\theta}$ intermediate state. From there, we see that the second photon has a higher probability of being emitted in a direction where $\cos\theta^{\prime}=\cos\theta$ to reach the |$\Ket{1,-1}$ ground state [19].
It is customary in studies of superradiance to set the initial conditions such that the ensemble of atoms is initially in a state of complete inversion (i.e., all atoms are in the excited state |$\Ket{b}$, as done in the previous paragraph [20]). However, in later sections we will be interested in situations where the system is in thermal equilibrium where all the states for a given $m$ value are equally populated. We will thus, from now on, focus on the transition rates between the $m=0$ and $m=-1$ states given in equations (13)-(14). To simplify our discussion we set $\cos\theta=0$ and integrate these transition rates over $d\Omega^{\prime}$ to find
with
[…]
$\left.+\left[1-3\left(\bm{\epsilon}_{\mu}\cdot\bm{\epsilon}_{z}\right)^{2}\right]\left[\frac{\cos\left(kz_{0}\right)}{\left(kz_{0}\right)^{2}}-\frac{\sin\left(kz_{0}\right)}{\left(kz_{0}\right)^{3}}\right]\right\}$ (17)
and
$\Gamma=\frac{\mu_{0}\omega_{0}^{3}\mu^{2}}{3\pi\hbar c^{3}}$ (18)
We consider initial conditions where the states |$\Ket{1,0}_{\theta}$ and |$\Ket{0,0}_{\theta}$ have equal probabilities of occupation (of $1/2$) but with uncorrelated probability amplitude coefficients. Under these conditions, the radiation intensity can be evaluated with
It is instructive to look at the characteristics of the system at opposite limits of separation between the two atoms. More precisely, in the limit $kz_{0}\rightarrow\infty$ where $F\left(kz_{0}\right)\rightarrow 0$ we find
@@
which is exactly the intensity from a single independent atom. In other words, when separated by an infinite distance the two entangled atoms will spontaneously emit one photon with a probability of unity over a time-scale $\Gamma^{-1}$ (when only one is in the excited state), just as an independent excited atom would.
On the other hand, in the idealized small sample limit $kz_{0}=0$ when $F\left(kz_{0}\right)=1$ the intensity becomes
@@
The small sample behavior is readily understood from equations (13)-(14)(or 15-17) since in this case coherence happens for all radiation modes, resulting in the doubling and cancellation of corresponding transition rates. The latter implies that |$\Ket{0,0}_{\theta}$ is a dark state and that its initial occupation probability is preserved at all times. This is where energy trapping originates. On the other hand, the behavior for $kz_{0}\rightarrow\infty$ is a global characteristic in the sense that it can only be discerned in the solid-angle-integrated transition rate of equation (16) through the function $F\left(kz_{0}\right)$. Even at very large separations some radiation modes will be coherent, but when considering all modes the two atoms behave as if they were independent.
@@
This last observation is important since external factors can sometimes render specific radiation modes dominant over the rest. For example, elongated (e.g., cylindrical) geometries are often considered in studies of superradiance since in such cases radiation along the symmetry axis is favored [19, 21, 20, 6]. Similarly, a gas in the interstellar medium initially hosting a population inversion will be more likely to radiate in a direction where velocity coherence is accentuated, leading to a transient superradiance response [6, 14]. Such considerations will be central to the material presented in Sec. IV.

### v2

# Quantum coherence and the invisible Universe: Subradiance as a dark matter mechanism

@@ Abstract
We investigate the role of quantum entanglement and coherence in suppressing radiation and explore its implications for dark matter. Using Dicke’s framework, we demonstrate that entangled states in a gas at thermal equilibrium can lead to subradiance, trapping energy in dark quantum states and reducing radiation intensity. Applying this to the 21 cm line in dark matter halos, we find that quantum coherence renders the gas effectively dark. Moreover, entanglement results in a vanishing collision cross-section, consistent with the collisionless nature of dark matter observed in systems like the bullet cluster. We also show that absorption of incident radiation can exceed levels predicted by Beer’s law, which holds only in the absence of coherence. These findings suggest that quantum entanglement and coherence may explain the non-luminous behavior of matter in dark matter halos, offering a novel perspective on dark matter and advancing the understanding of astrophysical radiative processes.
keywords:
entanglement, coherence, subradiance, dark matter
1 Introduction
It has been known since Dicke’s original treatment of the spontaneous emission of coherent radiation [1] that the interaction between atoms composing a gas [^1] Although we focus on atoms in this paper, our analysis and discussions apply to molecular gases as well. through their common electromagnetic field leads to the formation of entangled quantum mechanical states, with significant implications for the spontaneous emission rates of photons emanating from the gas. As Dicke pointed out, atoms should not be considered as independent entities; instead, the entire gas must be treated as a single quantum mechanical system. While the material covered in his paper was broad, touching on several phenomena involving coherence and entanglement, much of the early work it inspired focused on superradiance.
@@ 1 Introduction
Dicke’s 1954 paper [1] also explicitly discussed another coherent radiation process, which later became known as subradiance. In contrast to superradiance, subradiance describes the suppression of photon emission compared to the typical spontaneous emission process. Dicke analyzed the trapping of energy in slow and dark quantum states and its consequences for the radiation intensity emanating from a gas. Subradiance was experimentally verified later than superradiance [15] and has only recently been observed in large clouds of cold atoms [16]. The latter experimental result is particularly relevant to us as we explore the possible existence of subradiant systems in astrophysical contexts.
@@
2 Results
2.1 Entangled quantum states and coherent behavior
In this section we revisit parts of Dicke’s work and focus on the trapping of energy within an atomic ensemble. We first focus on a two-atom system at resonance with Atom 1 at $-z_{0}/2$ and Atom 2 at $z_{0}/2$ on the $z$-axis. Given that single-atom transitions take place between the lower |$\Ket{a}$ and upper |$\Ket{b}$ levels of energies $-\hbar\omega_{0}/2$ and $\hbar\omega_{0}/2$, respectively, appropriate internal states for the two-atom problem consist of [1, 18] (see Appendices A and D)
|$\Ket{1,1}$ =|$=\Ket{bb}$ (1)
@@ 2.1 Entangled quantum states and coherent behavior
Following Dicke [1], the kets |$\Ket{r,m}$ are defined with the “cooperative” number $r$ and “inversion” number $m$, which are analogous to spin-1/2 quantum numbers, with $r\geq 0$ and $\left|m\right|\leq r$ (|$\Ket{r,m}_{\theta}=\Ket{r,m}$ when $\cos\theta=0$). Note that $k=2\pi/\lambda$ with $\lambda$ the radiation wavelength.
Given equations (1)-(4) it is straightforward to calculate the spontaneous emission transition rates per unit solid angle between the different states
@@
The signatures of entanglement and coherence are displayed in equations (7)-(8). That is, the enhancement of the transitions rates by a factor of ‘2’ is due to the entangled nature of the states given in equations (2) and (4), while the disappearance of the argument in the cosine and sine functions (i.e., when $\cos\theta^{\prime}=\cos\theta$) stems from coherence between the two emitted photons. We therefore find the existence of an angular correlation between successively emitted photons. More precisely, if the system starts in the |$\Ket{1,1}$ state, then after the emission of the first photon it will for sure end up in the symmetric |$\Ket{1,0}_{\theta}$ intermediate state. From there, we see that the second photon has a higher probability of being emitted in a direction where $\cos\theta^{\prime}=\cos\theta$ to reach the |$\Ket{1,-1}$ ground state [18].
We will soon be interested in situations where the system is in thermal equilibrium where all the states for a given $m$ value are initially equally populated. We thus focus on the transition rates between the $m=0$ and $m=-1$ states given in equations (7)-(8). We can obtain the radiation intensity by, among other things, integrating these transition rates over the solid angle. Setting $\cos\theta=0$, for simplicity, and initial conditions where the states |$\Ket{1,0}_{\theta}$ and |$\Ket{0,0}_{\theta}$ have equal but uncorrelated probabilities of occupation (of $1/2$), we can study the intensity in two opposite limits (see Appendix A.1).
For the infinite size sample $kz_{0}\rightarrow\infty$ we find
@@
which is exactly the intensity from a single independent atom, with $\Gamma$ is the single-atom free-space spontaneous emission rate (see equation 34).
On the other hand, in the idealized small sample limit $kz_{0}=0$ the intensity becomes
@@
The small sample behavior is readily understood from equations (7)-(8) since in this case coherence happens for all radiation modes, resulting in the doubling and cancellation of the corresponding transition rates. The latter implies that |$\Ket{0,0}_{\theta}$ is a dark state and that its initial occupation probability is preserved at all times. This is where energy trapping originates. On the other hand, the behavior for $kz_{0}\rightarrow\infty$ is a global characteristic in the sense that it can only be discerned in the solid-angle-integrated transition rate. Even at very large separations some radiation modes will be coherent, but when considering all modes the two atoms behave as if they were independent (see equation 32).
@@
This last observation is important since external factors can sometimes render specific radiation modes dominant over the rest. For example, elongated (e.g., cylindrical) geometries are often considered in studies of superradiance since in such cases radiation along the symmetry axis is favored [18, 19, 20, 5]. Similarly, a gas in the interstellar medium initially hosting a population inversion will be more likely to radiate in a direction where velocity coherence is accentuated, leading to a transient superradiance response [5, 13].